OpenStudy (anonymous):

[2.06] Solve the equation 2/5 (x − 4) = 2x for x. −8 -1/2 1 −1

OpenStudy (anonymous):

First, divde both sides by 2/5 which is the same thing as multiplying by 5/2 2/5 (x − 4) = 2x x - 4 = 5x -4 = 4x x = -1

OpenStudy (anonymous):

-1